An inflation-indexed Treasury bond has a par value of $1,000 and a coupon rate of 2 percent.

An investor purchases this bond and holds it for one year. During the year, the consumer price

index increases by .25 percent every six months. What are the total interest payments the

investor will receive during the year
Answer #1

With Inflation indexed bonds, the value of bond gets adjusted for the inflation

Inflation adjusted value of bond after half year = 1000 * (1 + 0.25%)

= 1002.50

Coupon paid at half year = 1002.50 * 2%/2 = 10.025

Inflation adjusted value of bond after half year = 1002.50 * (1+0.25%) = 1005.06

Coupon paid at year 1 = 1005.06 * 2%/2 = 10.05